OBSERVATION OF NEW SHARP TRANSITIONS IN PARAHYDROGEN CRYSTALS DOPED WITH METHANE

\underline{C. MICHAEL LINDSAY} and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA; TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan.

In the course of studying ionized methane/hydrogen clusters in solid parahydrogen crystals, we accidentally stumbled upon a new set of infrared transitions in the crystal, apparently induced by the methane in the crystal (before ionization). Over 80 sharp absorption features (Δν ∼ 100 MHz) spanning 2 cm^{−1} were observed in the region between hydrogen’s Q_{1}(0) and Q_{1}(1) vibrational transitions. The spectroscopy of parahydrogen crystals in this region is well understood, and it was a complete surprise to observe the appearance of new transitions. In this talk we present the data and discuss our current progress in the understanding of this very unexpected spectrum.